OpenStudy (colcaps):

What are the length of the legs of a right triangle ABC if the legs are equal and the hypotenuse is 6 units? quare root of 2 ? square root of 3? square root of 5? square root of 6?

OpenStudy (colcaps):

i need help

OpenStudy (nubeer):

6^2 = x^2 + x^2 36 = 2x^2 x^2 = 18 x = \[3\sqrt{2}\] wel answer should be soemthing like this but it's not in option.. let me see again

OpenStudy (colcaps):

can you simplify it?

OpenStudy (nubeer):

hmm ok let me explain again... it says 2 sides are same and hyp is 6 means base and height are equal.. so we can say the length of one side is. length = x , height = x , hyp =6 now apply pythagoras hyp^2= h^2 + b^2 6^2 = x^2 + x^2 36 = 2x^2 x^2 = 18 x = 3√2